# Nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes Equation

### v1

@@ Abstract
The nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes equations is one of the central open problems in mathematical fluid dynamics. In this paper, we provide, to our knowledge, the first rigorous computer-assisted proof demonstrating such nonuniqueness. Inspired by earlier works [30, 1, 25], we construct a Leray–Hopf solution in the self-similar setting and then establish the existence of a second solution by analyzing the stability of the linearized operator around this profile, showing that it corresponds to an unstable perturbation. To achieve this, we develop an innovative numerical method that computes candidate solutions with high precision and propose a framework for rigorously establishing exact solutions in a neighborhood of these candidates. A key step is to decompose the linearized operator into a coercive part plus a compact perturbation, which is further approximated by a finite-rank operator up to a small error. The invertibility of the linearized operator restricted to the image of this finite-rank approximation is then rigorously verified using computer-assisted proofs. This certifies the existence of an unstable eigenpair and, consequently, yields a second solution–indeed, infinitely many Leray–Hopf solutions.
@@ 1 Introduction
There exist infinitely many distinct suitable Leray–Hopf solutions to the Navier–Stokes equation (1.1) on $\mathbb{R}^{3}\times[0,1]$ with the same divergence-free initial condition $u_{\operatorname{in}}=u_{\operatorname{loc}}$ of compact support, with $u_{\operatorname{loc}}\in L^{q}$ for any $q<3$ (in particular $u_{\operatorname{in}}\in L^{2}$). The solutions are $L^{s}([0,1];L^{q}(\mathbb{R}^{3})),$ for any $q\geqslant 2$, $\frac{3}{q}+\frac{2}{s}>1$. Namely, they just miss the Prodi–Serrin condition that guarantees a unique solution, see (1.3).
@@
In this paper, we adopt a different strategy by developing a novel computational method and an innovative computer-assisted strategy to rigorously verify the existence of a self-similar profile that solves the unforced Navier–Stokes equation exactly with an unstable eigenvalue. We then use the tools developed in [1, 30] to establish the nonuniqueness of Leray–Hopf solutions.
@@ 1.1 Leray–Hopf solution and strong-weak uniqueness
The critical case of $q=3$ was established by Escauriaza–Seregin–Sveark [23]. Under this condition, the solution exhibits higher regularity [42, 43, 23],
@@ 1.3 Eigenvalue of the linearized operator and nonuniqueness
A subtle point is the regularity and decay of the exact profile $\widetilde{U}:=\overline{U}+U$. From Proposition 1 we only know $U\in H^{1}$ with $\|U\|_{H^{1}}\leqslant\varepsilon^{U}$ (via Schauder), which does not by itself identify $\widetilde{U}$ as a Leray–Hopf profile. To bridge this, we introduce a smooth radial cutoff at scale $R\gg 1$ and write $\widetilde{U}=U_{\mathrm{near}}+U_{\mathrm{far}}$, where $U_{\mathrm{far}}$ is divergence free, smooth, and constructed to match the exterior asymptotics of $\overline{U}$ (hence the same far–field boundary condition and $|\xi|^{-1}$ decay). In these variables, $U_{\mathrm{near}}$ solves the stationary self-similar Navier–Stokes system with variable coefficients determined by $U_{\mathrm{far}}$; these coefficients are smooth, $C^{k}$–bounded, and decay polynomially, and the corresponding forcing is smooth and decays like $O(|\xi|^{-3})$ by our far-field analysis. Because neither the coefficients nor the forcing are compactly supported, standard local estimates are insufficient. We therefore combine local elliptic bootstrapping (for variable, decaying coefficients) with the coercivity of the linearized operator to obtain $U_{\mathrm{near}}\in H^{m}$ for all $m$, together with weighted energy bounds
@@
We combine the localization idea of Jia-Sverak [30] with the fixed-point framework of Albritton-Brue-Colombo [1]. Writing the solution as $u=\widetilde{u}+u^{\mathrm{cut}}+u^{\mathrm{cor}}$, we first localize the scale-invariant initial data by a Bogovskii decomposition $u_{\mathrm{in}}=u_{\operatorname{loc}}+w$, where $u_{\operatorname{loc}}$ is compactly supported and divergence free, and then set $u^{\mathrm{cut}}=-e^{t\Delta}w$ to peel off the far-field tail. Unlike [1], our self-similar profile $\widetilde{U}$ is not compactly supported; the new ingredient is to control the resulting far-field interaction via heat smoothing, which makes the associated linear term small when the localization radius $R$ is large. Passing to similarity variables, we solve for the correction $U^{\mathrm{cor}}$ as a fixed point of a Duhamel map generated by the semigroup $e^{\tau\mathcal{L}_{\widetilde{U}}}$. A key idea is to use a Riesz projection to split the dynamics into a stable part and a unstable part as in [30]; see also [11, 10]. We will show that the map is a contraction. The finitely many unstable modes will serve as free parameters at $\tau=0$. Different choices of these unstable coordinates yield distinct solutions on $(-\infty,0]$, producing infinitely many Leray–Hopf solutions with the same $L^{2}$ initial data $u_{\operatorname{loc}}$. In this way we retain the advantages of both approaches–localization from [30] and the fixed-point mechanism of [1]–while handling a non-compactly supported self-similar background.
In our construction of the Leray–Hopf solution $\overline{U}$ and the eigenfunction $\overline{v}$, we impose even symmetry in $z$ and odd symmetry in $z$ for $\overline{v}$. The odd parity of $\overline{v}$ breaks the base even symmetry of $\overline{U}$; thus the Leray–Hopf branch bifurcating in the $\overline{v}$-direction is symmetry-breaking, which was also observed by the previous work [25, 1].
@@ 1.4 Computer-assisted proofs
We would like to point out that our Navier–Stokes nonuniqueness verification is far lighter than that for the 3D Euler singularity program [12, 14, 13], because viscosity yields genuine coercivity. In particular, the linearized operators enjoy relatively large coercivity coefficients in $L^{2}$: about $0.054$ for the $\overline{U}$–equation and about $0.16$ for the $\overline{v}$–equation (see Section 4.5). This allows us to work in standard $L^{2}/H^{1}$ norms without singular weights and to close the a posteriori bounds using only the $L^{2}$ residuals of $\overline{U}$ and $\overline{v}$ together with a few pointwise envelopes. The discrete operators we assemble are modest (roughly $1200\times 1200$), and the finite-rank verification reduces to a small, static generalized eigenproblems of size $25\times 25$. Thanks to the strong damping reflected by the above coercivity and the effective approximation property of our generalized eigenfunctions, retaining $25$ modes suffices to close the estimates via the Schauder fixed point theorem.
Moreover, we use a trigonometric tensor-product representation in $(\beta,\theta)$ for $\overline{U}$ and $\overline{v}$ after the change of variables $r=\tan\beta$ with a relatively small number of modes ($N=600$ in the $\beta$-direction and $N=300$ in the $\theta$-direction). This representation enables us to perform analytic $L^{2}$ evaluation of the residuals $E^{U}$ and $E^{v}$ (no round-off accumulation) and also makes it easy for us to obtain relatively sharp upper bounds for $\|\overline{U}\|_{L^{\infty}}$ and $\|\overline{v}\|_{L^{\infty}}$ as well as their gradients. The only delicate numerical step is to evaluate the $Q$-inner products for the generalized eigenfunctions which involve Bessel functions along the $\beta$-direction. We need to approximate the inner products using a 4th order quadrature and supply a rigorous $W^{4,\infty}$ error bound on the integrand, see Section 7 for more discussion. Altogether, viscous coercivity, low-dimensional spectral checks, and analytic trigonometric calculus make the present verification substantially simpler than in the Euler setting.
We remark that a candidate solution was found in [25], but this approximate solution seems to be too large to apply our computer-assisted proof, in the sense that the apprximation of the associated compact part of the linearized operator by a finite rank operator would require a large number of modes to reduce the residual error small enough for us to close the bootstrap argument, making the computer verification too expensive; see Section 4.2 for details.
@@ 1.6 Literature review
One of the earliest examples of nonuniqueness to (1.1) dates back to Ladyzhenskaya [32], in a time-varying domain with force. Similar to the program of Jia, Sverak, and Guillod, Bressan, Murray, and Shen [5, 4] proposed some suggestive numerical evidence for nonuniqueness to the Euler equations without forcing. In addition to the work using self-similar solutions, there is another important line of work leveraging convex integration, a powerful tool introduced to fluid dynamics that resulted in the proof of the Onsager conjecture by Isett [28] and Buckmaster–De-Lellis–Szekelyhidi–Vicol [6]. In particular, Buckmaster–Vicol [7] established that for any $L^{2}$ initial data, distributional solutions to (1.1) are nonunique, and Cheskidov–Luo [18, 19] established nonuniqueness in $L_{t}^{2-}L_{x}^{\infty}$ and $L_{t}^{\infty}L_{x}^{2-}$ for dimension $d\geqslant 2$. However, solutions obtained from convex integration schemes are very far from Leray regularity $H^{1}_{x}$. Very recently, Coiculescu–Palasek established nonuniqueness in a scaling critical space $BMO^{-1}$ [20] where global well-posedness for small initial data holds [31], using techniques from a dyadic model introduced by Palasek [41].
@@ 2.1 Regularity of the profile
Sketch of proof for Proposition 3. $U_{\operatorname{far}}$ is smooth, so we only need to show the smoothness of $U_{\operatorname{near}}$. We defer the full proof to Appendix G and outline the main steps of coercive estimates here.
@@
$\|\rho_{\epsilon}*\nabla^{k-1}\Pi(\widetilde{U}\cdot\nabla\widetilde{U})\|_{L^{2}}\leqslant C_{k}(1+\|U_{\mathrm{near}}\|^{2}_{H^{k}}+\|u^{\epsilon}\|_{L^{\infty}}\|U_{\mathrm{near}}\|_{H^{k}})\leqslant C_{k}(1+\|U_{\mathrm{near}}\|^{2}_{H^{k}}+\|u^{\epsilon}\|_{H^{k+1}}^{3/4}\|U_{\mathrm{near}}\|_{H^{k}}^{5/4}),$
@@
Sketch of Proof for Proposition 4. Since $U_{\operatorname{far}}$ satisfies the desirable far field condition (2.5), we only need to establish $U_{\operatorname{near}}$ has decay $O(|\xi|^{-3/2})$ in the far field. We will establish that $|\xi|U_{\operatorname{near}},|\xi|^{2}\nabla^{2}U_{\operatorname{near}}\in L^{2}$, and use a dyadic Sobolev embedding to conclude the far field decay. We defer the full proof to Appendix G and outline the weighted estimates and dyadic rescaling.
@@
$\phi_{L}(\xi):=\phi\left(|\xi|/L\right),\qquad\phi(r)=r^{2}\ (0\leqslant r\leqslant 1),\quad\phi(r)=4\ (r\geqslant 2).$
@@
where we use the fact by (2.5) and (2.6) that $|\xi|\nabla U_{\operatorname{far}},\widetilde{U}\in L^{4}\cap L^{\infty}$, and the uniform bounds on derivatives of $\phi_{L}$. For the pressure, we use the Riesz transform to represent $\Delta P=-\partial_{i}\partial_{j}(\widetilde{U}_{i}\widetilde{U}_{j})$ and Calderón–Zygmund theory give $\|P\|_{L^{2}}\lesssim\|\widetilde{U}\|^{2}_{L^{4}}\lesssim 1$, hence
$|\langle\nabla P,\,\phi_{L}U_{\mathrm{near}}\rangle|=|\langle P,\,\nabla\phi_{L}\cdot U_{\mathrm{near}}\rangle|\lesssim\sqrt{\langle U_{\operatorname{near}},\phi_{L}U_{\operatorname{near}}\rangle}.$
The forcing term has the same bound by the decay of $F$. Collecting the estimates yields a uniform bound on $\langle U_{\operatorname{near}},\phi_{L}U_{\operatorname{near}}\rangle+\langle\nabla U_{\operatorname{near}},\phi_{L}\nabla U_{\operatorname{near}}\rangle$, and we can take the limit $L\rightarrow\infty$ to conclude
@@
Dyadic rescaling and decay. Fix $R>1$ and the annulus $A_{R}=\{\xi\in\mathbb{R}^{3}:\ R<|\xi|<2R\}$. Define $v(y):=U_{\mathrm{near}}(Ry)$ on $A_{1}=\{y:\ 1<|y|<2\}$. By the local Sobolev embedding $H^{2}(A_{1})\subset L^{\infty}(A_{1})$,

### v2

@@ Abstract
The nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes equations is one of the central open problems in mathematical fluid dynamics. In this paper, we provide, to our knowledge, the first rigorous computer-assisted proof demonstrating such nonuniqueness. Inspired by earlier works [32, 1, 27], we construct a Leray–Hopf solution in the self-similar setting and then establish the existence of a second solution by analyzing the stability of the linearized operator around this profile, showing that it corresponds to an unstable perturbation. To achieve this, we develop an innovative numerical method that computes candidate solutions with high precision and propose a framework for rigorously establishing exact solutions in a neighborhood of these candidates, where the invertibility of the linearized operator is the key. Inspired by [15], we decompose the linearized operator into a coercive part plus a compact perturbation, which is further approximated by a finite-rank operator up to a small error. The invertibility of the linearized operator restricted to the image of this finite-rank approximation is then rigorously verified via construction of the inversion on its finite set of basis functions with residues incorporated properly. This certifies the existence of a profile and similarly an unstable eigenpair. Consequently, it yields a second solution–indeed, infinitely many Leray–Hopf solutions. The code is publicly available at: https://github.com/HouGroup2026/3d-navier-stokes-nonuniqueness
@@ 1 Introduction
There exist infinitely many distinct suitable Leray–Hopf solutions to the Navier–Stokes equation (1.1) on $\mathbb{R}^{3}\times[0,1]$ with the same divergence-free initial condition $u_{\operatorname{in}}=u_{\operatorname{loc}}$ of compact support, with $u_{\operatorname{loc}}\in C^{\infty}(\mathbb{R}^{3}\setminus\{0\})\cap L^{q}$ for any $q<3$. The solutions are $L^{s}([0,1];L^{q}(\mathbb{R}^{3})),$ for any $q\geqslant 2$, $\frac{3}{q}+\frac{2}{s}>1$, and smooth for positive times. They just miss the Prodi–Serrin condition that guarantees a unique solution, see (1.3).
@@
In this paper, we adopt a different strategy by developing a novel computational method and an innovative computer-assisted strategy to rigorously verify the existence of a self-similar profile that solves the unforced Navier–Stokes equation exactly with an unstable eigenvalue. We then use the tools developed in [32] to establish the nonuniqueness of Leray–Hopf solutions, while reducing the spectral assumption simply to an unstable eigenvalue inspired by [1].
@@ 1.1 Leray–Hopf solution and strong-weak uniqueness
The critical case of $q=3$ was established by Escauriaza–Seregin–Sverak [25]. Under this condition, the solution exhibits higher regularity [44, 45, 25],
@@ 1.3 Eigenvalue of the linearized operator and nonuniqueness
A subtle point is the regularity and decay of the exact profile $\widetilde{U}:=\overline{U}+U$. From Proposition 1 we only know $U\in H^{1}$ with $\|U\|_{H^{1}}\lesssim\varepsilon^{U}$, which does not by itself identify $\widetilde{U}$ as a Leray–Hopf profile. To bridge this, we introduce a smooth radial cutoff and write $\widetilde{U}=U_{\mathrm{near}}+U_{\mathrm{far}}$, where $U_{\mathrm{far}}$ is divergence free, smooth, and constructed to match the exterior asymptotics of $\overline{U}$. In these variables, $U_{\mathrm{near}}$ solves the stationary self-similar Navier–Stokes system with variable coefficients determined by $U_{\mathrm{far}}$; these coefficients are smooth, $C^{k}$–bounded, and decay polynomially, and the corresponding forcing is smooth and decays like $O(|\xi|^{-3})$ by our far-field analysis. We therefore combine local elliptic bootstrapping using the coercivity of the linearized operator to obtain $U_{\mathrm{near}}\in H^{m}$ for all $m$, together with weighted energy bounds
@@
We combine the localization idea of Jia–Sverak [32] with the fixed-point framework of Jia–Sverak [32] and Albritton–Brue–Colombo [1]. The key innovation compared to [32] is that we perform localization and nonuniqueness arguments at the same time, bypassing the sophisticated spectral assumptions therein. Writing the solution as $u=\widetilde{u}+u^{\mathrm{cut}}+u^{\mathrm{cor}}$, we first localize the scale-invariant initial data by a Bogovskii decomposition $u_{\mathrm{in}}=u_{\operatorname{loc}}+w$, where $u_{\operatorname{loc}}$ is compactly supported and divergence free, and then set $u^{\mathrm{cut}}=-e^{t\Delta}w$ to cut off the far-field tail. Unlike [1], our self-similar profile $\widetilde{U}$ is not compactly supported; the new ingredient is to control the resulting far-field interaction via heat smoothing, which makes the associated linear term small when the localization radius $R$ is large. Passing to similarity variables, we solve for the correction $U^{\mathrm{cor}}$ as a fixed point of a contraction Duhamel map generated by the semigroup $e^{\tau\mathcal{L}_{\widetilde{U}}}$. A key idea is to use a Riesz projection to split the dynamics into a stable part and an unstable part as in [32]; see also [12, 18]. The finitely many unstable modes will serve as free parameters at $\tau=0$. Different choices of these unstable coordinates yield distinct solutions on $(-\infty,0]$, producing infinitely many Leray–Hopf solutions with the same $L^{2}$ initial data $u_{\operatorname{loc}}$.
In our construction of the Leray–Hopf solution $\overline{U}$ and the eigenfunction $\overline{v}$, we impose even symmetry in $z$ for $\overline{U}$ and odd symmetry in $z$ for $\overline{v}$. The odd parity of $\overline{v}$ breaks the base even symmetry of $\overline{U}$; thus the Leray–Hopf branch bifurcating in the $\overline{v}$-direction is symmetry-breaking, which was also observed by the previous work [27, 1].
@@ 1.4 Computer-assisted proofs
We would like to point out that our Navier–Stokes nonuniqueness verification is far lighter than that for the 3D Euler singularity program [15, 17, 16], because viscosity yields genuine coercivity. In particular, the linearized operators enjoy relatively large coercivity coefficients in $L^{2}$: about $0.051$ for the $\overline{U}$–equation and about $0.164$ for the $\overline{v}$–equation (see Section 4.5). This allows us to work in standard $L^{2}/H^{1}$ norms without singular weights and to close the a posteriori bounds using only the $L^{2}$ residuals of $\overline{U}$ and $\overline{v}$ together with a few pointwise envelopes of $L^{\infty}$ estimates. The discrete operators we assemble are modest (roughly $1200\times 1200$), and the finite-rank verification reduces to eigenproblems of small sizes $25\times 25$. Thanks to the strong damping reflected by the above coercivity and the effective approximation property of our generalized eigenfunctions, retaining $25$ modes suffices to close the estimates via the Schauder fixed point theorem.
Moreover, we use a trigonometric tensor-product representation in $(\beta,\theta)$ for $\overline{U}$ and $\overline{v}$ after the change of variables $r=\tan\beta$ with a relatively small number of modes ($N=600$ in the $\beta$-direction and $N=300$ in the $\theta$-direction). This representation enables us to perform analytic $L^{2}$ evaluation of the residuals $E^{U}$ and $E^{v}$ and also makes it easy for us to obtain relatively sharp upper bounds for $\|\overline{U}\|_{L^{\infty}}$ and $\|\overline{v}\|_{L^{\infty}}$ as well as their gradients. The only delicate numerical step is to evaluate the $Q$-inner products for the generalized eigenfunctions, which involve Bessel functions along the $\beta$-direction. We need to approximate the inner products using a 8th order quadrature and supply a rigorous $W^{8,\infty}$ error bound on the integrand, see Section 7 for more discussion. Altogether, viscous coercivity, low-dimensional spectral checks, and analytic trigonometric calculus make the present verification substantially simpler than that in the Euler setting.
We remark that a candidate solution was found in [27], but this approximate solution seems to be too large to apply our computer-assisted proof, in the sense that the approximation of the associated compact part of the linearized operator by a finite rank operator would require a large number of modes to reduce the residual error small enough for us to close the bootstrap argument, making the computer verification too expensive; see Section 4.2 for details.
@@ 1.6 Literature review
One of the earliest examples of nonuniqueness to (1.1) dates back to Ladyzhenskaya [34], in a time-varying domain with force. Similar to the program of Jia, Sverak, and Guillod, Bressan, Murray, and Shen [6, 5] proposed some suggestive numerical evidence for nonuniqueness to the 2D Euler equations without forcing. In addition to the work using self-similar solutions, there is another important line of work leveraging convex integration, a powerful tool introduced to fluid dynamics that resulted in the proof of the Onsager conjecture by Isett [30] and Buckmaster–De-Lellis–Szekelyhidi–Vicol [7]. In particular, Buckmaster–Vicol [8] established that for any $L^{2}$ initial data, distributional solutions to (1.1) are nonunique, and Cheskidov–Luo [20, 21] established nonuniqueness in $L_{t}^{2-}L_{x}^{\infty}$ and $L_{t}^{\infty}L_{x}^{2-}$ for dimension $d\geqslant 2$. However, solutions obtained from convex integration schemes are very far from Leray regularity $H^{1}_{x}$. Very recently, Coiculescu–Palasek established nonuniqueness in a scaling critical space $BMO^{-1}$ [22] where global well-posedness for small initial data holds [33], using techniques from a dyadic model introduced by Palasek [43]. Later on, Cheskidov–Dai–Palasek [19] established nonuniqueness in the same space along with instantaneous blowup.
@@ 2.1 Regularity of the profile
Sketch of proof for Proposition 3. $U_{\operatorname{far}}$ is smooth, so we only need to show the smoothness of $U_{\operatorname{near}}$. We defer the full proof to Appendix F and outline the main steps of coercive estimates here.
@@
$\|\rho_{\epsilon}*\nabla^{k-1}\Pi(\widetilde{U}\cdot\nabla\widetilde{U})\|_{L^{2}}\leqslant C_{k}(1+\|U_{\mathrm{near}}\|^{2}_{H^{k}}+\|u^{\epsilon}\|_{L^{\infty}}\|U_{\mathrm{near}}\|_{H^{k}})\leqslant C_{k}(1+\|U_{\mathrm{near}}\|^{2}_{H^{k}}+\|u^{\epsilon}\|_{H^{k+1}}^{3/4}\|U_{\mathrm{near}}\|_{H^{k}}^{5/4}).$
@@
Sketch of Proof for Proposition 4. Since $U_{\operatorname{far}}$ satisfies the desirable far field condition (2.5), we only need to establish $U_{\operatorname{near}}$ has decay $O(|\xi|^{-3/2})$ in the far field. We will establish that $|\xi|U_{\operatorname{near}},|\xi|^{2}\nabla^{2}U_{\operatorname{near}}\in L^{2}$, and use a dyadic Sobolev embedding to conclude the far field decay. We defer the full proof to Appendix F and outline the weighted estimates and dyadic rescaling.
@@
$\phi_{L}(\xi):=\phi\left(|\xi|/L\right),\qquad\phi(r)=r^{2}\ (0\leqslant r\leqslant 1),\quad\phi(r)=4\ (r\geqslant 2),$
@@
where we use the fact by (2.5) and (2.6) that $|\xi|\nabla U_{\operatorname{far}},\widetilde{U}\in L^{4}\cap L^{\infty}$, and the uniform bounds on derivatives of $\phi_{L}$. For the pressure, we use the Riesz transform to represent $\Delta P=-\partial_{i}\partial_{j}(\widetilde{U}_{i}\widetilde{U}_{j})$ and the Calderón–Zygmund theory gives $\|P\|_{L^{2}}\lesssim\|\widetilde{U}\|^{2}_{L^{4}}\lesssim 1$, hence we obtain
$|\langle\nabla P,\,\phi_{L}U_{\mathrm{near}}\rangle|=|\langle P,\,\nabla\phi_{L}\cdot U_{\mathrm{near}}\rangle|\lesssim\sqrt{\langle U_{\operatorname{near}},\phi_{L}U_{\operatorname{near}}\rangle},$
where we have used the fact that $|\nabla\phi_{L}|^{2}\lesssim\phi_{L}$. The forcing term has the same bound by the decay of $F$. Collecting the estimates yields a uniform bound on $\langle U_{\operatorname{near}},\phi_{L}U_{\operatorname{near}}\rangle+\langle\nabla U_{\operatorname{near}},\phi_{L}\nabla U_{\operatorname{near}}\rangle$, and we can take the limit $L\rightarrow\infty$ to conclude
@@
Dyadic rescaling and decay. Fix $R>1$ and the annulus $A_{R}=\{\xi\in\mathbb{R}^{3}:\ R<|\xi|<2R\}$. Define $v(y):=U_{\mathrm{near}}(Ry)$ on $A_{1}=\{y:\ 1<|y|<2\}$. The local Sobolev embedding $H^{2}(A_{1})\subset L^{\infty}(A_{1})$ implies